Nutrition Facts Serving Size 1/4 Cup (30g) Servings Per Container About 10 Amount Per Serving Calories 150 Calories from Fat 70 % Daily Value* 12% 396 Total Fat 8g Saturated Fat 0 Trans Fat 0g Polyunsaturated Fat 2.5g Monounsaturated Fat Cholesterol Omg Sodium Total Carbohydrate 14g 0% 1% 5% 10% Dietary Fiber 3g Sugars Protein 6g 7% Vitamin A 0% Vitamin C 0% Calcium 4% . Iron 6% Percent Daily Values are based on a 2,000 calorie diet. Your daily values may be higher or lower depending on your calorie needs. What is the serving size? How many servings are in the entire container? How many calories would you consume if you consumed 3.5 servings of this product? How many grams of total fat are in each serving? In one serving, how many calories are protein?
Answer #1

1. The serving size of the food item is 30grams

2. The number of servings in the entire container is 10 .

3. The amount of calories gained by consuming 3.5 serving of the product is :-

The amount of calories per serving is 150 calories,then 3.5 serving will provide = 3.5 × 150 = 525 calories .

4. The amount of total fat in each serving is 8grams.

5. In one serving of this food product ,the amount of protein is 6grams and 1gram of protein provides 4kcal .

Then 6grams will provide =6×4 = 24kcal .